OpenStudy (anonymous):

The school cafeteria sells two kinds of wraps: vegetarian and chicken. The vegetarian wrap costs $1.00 and the chicken wrap costs $1.80. Today they made $98.80 from the 70 wraps sold. How many of the wraps sold were vegetarian?

OpenStudy (anonymous):

V = 34. 34 vegetarian wraps sold Josh, this is the History section, not the Math section :) .

OpenStudy (anonymous):

LOL no one was responding over there tho but thanx

OpenStudy (anonymous):

Don't you have any notes to help you? :O

OpenStudy (anonymous):

i dont take notes lol

OpenStudy (anonymous):

>.> Start taking notes! It'll help you *a lot* in the long run :)

OpenStudy (anonymous):

Here's a helpful website for systems of equations: http://www.sosmath.com/soe/SE/SE.html Here's another in case the first one doesn't help enough: http://www.mathwarehouse.com/algebra/linear_equation/systems-of-equation/index.php

OpenStudy (anonymous):

Oh, if you could please, a medal :)

OpenStudy (anonymous):

lol ok. and thanks

OpenStudy (anonymous):

No problem :)

OpenStudy (texaschic101):

v stands for vegetarian wraps c stands for chicken wraps v + c = 70 ---> v = 70 - c 1.8c + 1v = 98.80 now sub 70 - c in for v in the second equation... 1.8c + 1(70 - c) = 98.80 (distribute the 1 through the parenthesis) 1.8c + 70 - c = 98.80 (combine like terms) .8c + 70 = 98.80 (subtract 70 from both sides) .8c + 70 - 70 = 98.80 - 70 .8c = 28.8 (divide by .8 to get c by itself) .8c/.8 = 28.8/.8 c = 36 now sub 36 in for c in the first equation... v + c = 70 v + 36 = 70 v = 70 - 36 v = 34 check... 1.8c + 1v = 98.80 1.8(36) + 1(34) = 98.80 64.8 + 34 = 98.80 98.80 = 98.80 (correct) ANSWER : vegetarian wraps sold were 34 at $ 1 each = $ 34.00 chicken wraps sold were 36 at $ 1.80 each = $ 64.80. total : 70 wraps at a total of $ 98.80